The sum of the ages of the two brothers is 20.In five years, one is twice the other. Find the present ages of the two brothers
Answer by ptfile(81)

x = the age of the older brother
y = the age of the younger brother
The sum of the ages of the two brothers is 20.
x+y=20
y=20-x
In five years, one is twice the other.
x+5=2(y+5)
x+5=2(20-x+5)
x+5=40-2x+10
x+5=50-2x
3x=45
x=15
y=5
The older brother's age is 15 years old while the younger brother's age is 5 years old.